OpenStudy (anonymous):

How many grams of nitrogen are in a flask with a volume of 250 mL at a pressure of 1500mm of Hg and a temperature of 27 C?

OpenStudy (anonymous):

Some how I got 426515.6, but i am almost definite that this is incorrect.

OpenStudy (perl):

did you use the three gas laws

OpenStudy (anonymous):

I plugged in 1500(250)=n(.08206)(300) And then I tried to convert the moles into grams

OpenStudy (perl):

we have three gas laws, where we assume ideal gases 1. PV = k (temperature held constant) 2. V/T = k (pressure held constant) 3. P/T= k (volume held constant) The combined gas law is PV = kT Avagadro's gas law is PV = nRT http://chemistry.bd.psu.edu/jircitano/gases.html

OpenStudy (perl):

so lets use avagadros gas law, we need to use kelvin , k = 273.15 + C

OpenStudy (perl):

to use avagadro's gas law, R = .08206, V is in liters, P is in atm's and T is in kelvins , n = gas molar mass

OpenStudy (perl):

and nitrogen is diatomic , dont forget

OpenStudy (perl):

first convert 1500mm of Hg into atm

OpenStudy (perl):

google says 1500 mm Hg =1.97368421 atm

OpenStudy (perl):

PV = nRT (1.97368421 atm)(.250 L) = n * .08206 * (27 + 273.15 Kelvins)

OpenStudy (perl):

solve for n, that gives you number of moles of N2

OpenStudy (perl):

PV = nRT (1.97368421 atm)(.250 L) = n * (.08206 L-atm / mol-K) * (27 + 273.15 K) notice that the units cancel n = (1.97368421 * .250 / (.08206 * 300.15) moles of N2

OpenStudy (perl):

n = .02 moles roughly now N2 is 28 grams / mole

OpenStudy (perl):

(.02 moles of N2 ) *( 28 g N2 ) / ( mole of N2) = 0.56 grams of N2

OpenStudy (perl):

is that correct?

OpenStudy (perl):

also you can use wolfram to find the answer http://www.wolframalpha.com/input/?i=mass+of+nitrogen+gas+250+mL+at+1500mm+of+Hg+at+temperature+of+27+Celsius